# EEG Emotion Copilot: Pruning LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

### v1

@@ III-D Model Deployment in EEG Emotion Copilot
Given the recent release of various model deployment tools, such as llama.cpp, which can quantify the model based on the specific scenario and convert it into gguf format, it can be directly utilized through ollama, llmstudio, or other tools. This process ultimately results in the model deployment part illustrated in Figure 1. Consequently, we will not delve into the details of model deployment and will proceed directly to the model application section.
III-E Assisted Electronic Medical Records
Different models need to be trained for various tasks. For emotion recognition, diagnosis, and treatment plan design, it is often sufficient to embed demographic data of different cases under various emotions along with the corresponding compressed EEG signals in the training data. In downstream applications, the collected participant data can be input into the fine-tuned model in a specific format to obtain the corresponding diagnosis results and treatment plans, as shown in Figure 5.
@@ IV-A Implementation Details
Dataset: We utilize the FACED dataset [24] to construct our training dataset. This dataset includes recordings from 123 subjects, each with 32 EEG channels, as they watched 28 video clips, each lasting 30 seconds. These clips encompass nine emotions: Anger, Disgust, Fear, Sadness, Neutral, Amusement, Inspiration, Joy, and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 and trained on the public dataset with a learning rate of 1e-5, incorporating L1 and L2 regularization terms. The training was conducted for 100 iterations. During the Lora fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRA_r to 8, LoRA_alpha to 32, and LoRA_dropout to 0.5. The fine-tuning process spanned 15 epochs, divided into 5 steps. Model pruning was executed on an A800 80G, while model training and fine-tuning were completed on a 4090 24G.
@@ IV-B Performance Validation of Varied EEG Signal Length
Regarding the compressed signal length, we first constructed the signal in the FACED data and the corresponding demography according to the Data Structure in Figure 1. We implemented two types of compression methods and an original control group. The question-answering process and the corresponding results are shown in Figure 6. Although the Euclidean distance and Cosine similarity of Ans1 and Ans2 are similar to the Baseline for the same question, Ans1 is more consistent with the Baseline than Ans2, as Ans2 differs from the Baseline in terms of emotion recognition. Therefore, with limited resources, it is often more effective to compress the signal to a fixed length using wavelet transforms rather than a segmented approach.
IV-C Model Retraining Strategy Selection After Pruning
We explored two sequential retraining strategies to restore model performance after applying torch pruning, as illustrated in Figure 7. It is reasonable to first train the model on public datasets and then on a specific dataset, as public datasets are rich in categories and provide a comprehensive feature representation for the model’s learning. This approach allows the model to leverage the diversity of public datasets to learn more generalized features initially. Consequently, when the model encounters a specific dataset, it can quickly adjust and optimize its weights to meet the specific task requirements.
Alternatively, we first trained on a specific dataset and then fine-tuned on a public dataset. This method allows the model to initially focus on the features of a specific task and subsequently improve its generalization ability and robustness through exposure to a public dataset. However, to save computing resources, training solely on a specific dataset may suffice.
We conducted an experiment for validation, as shown in Figure 8, dividing the training process into five steps, each with three epochs. In the first four steps, we used 250Hz data from 55 subjects in the FACED dataset. We observed that the model’s answers in step 3 were more satisfactory than in step 4, despite minimal differences in training_loss, indicating that excessive training may be unnecessary or even detrimental for lightweight models. In step 5, we used 1000Hz data from 68 subjects in the FACED dataset, with the pre-trained from step 4. The model trained on a dataset with a similar structure but different information density provided richer answers than in step 4, confirming that the model had learned the data structure and corresponding answers. This validates our hypothesis that the pruned model can be effectively retrained on a specific dataset to provide specific answers.
IV-D Assisted Electronic Medical Record Generation with Our Copilot
@@
Figure 9 illustrates an example of using our EEG Emotion Copilot to create an electronic medical record. Initially, the starting sentence of the prompt is provided, followed by the addition of demographic information of the subjects, including age, gender, and certain facial features. Finally, the preprocessed and compressed signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot. The EEG Emotion Copilot first analyzes the data to generate an answer, which is subsequently transformed into the record through format conversion.
Figure 6: Two demos of paired question and answer with signal lengths of 50 (model 1) and 500 (model 2) time points. The doctor first asks a question to the trained model 1 or model 2. For model 1, the wavelet transform compresses the original signal of each channel into 50 points as training data. For model 2, the wavelet transform initially compresses the original signal of each channel into 500 points, then divides the compressed channel signal into 10 segments. These ten sequential group data serve as training data for model 2. Ans1 and Ans2 are the answers from model 1 and model 2, respectively, while the baseline represents the real answer to the question.
Figure 7: Two types of retraining strategies. The first strategy involves initially training the model on a specific dataset tailored to the current task, followed by training on a public dataset. The second strategy begins with training the model on a public dataset, and subsequently fine-tuning it on a specific dataset.
Figure 8: Comparative experimental results of the model’s training process. The training is divided into five steps, each consisting of three epochs. The first four steps utilize 250Hz data from 55 subjects. Comparing the model’s answers in steps 3 and 4 reveals that the answer in step 3 is richer in content, with no significant difference in training loss between the two steps. This suggests that overtraining lightweight models may be redundant or even detrimental. In step 5, 1000Hz data from 68 subjects is used for training, employing the pre-trained model from step 4.
@@
V Discussion and Conclusion
In this paper, we present EEG Emotion Copilot, a lightweight and locally-running LLM designed for emotional recognition using EEG signals and the generation of corresponding treatment plans for assisted electronic medical records. Our approach addresses several key challenges in the fields of AC and BMIs.
Firstly, we developed a comprehensive human-based dataset that integrates public knowledge datasets and specialized emotion datasets, enabling the construction of a robust data structure to drive the development of LLMs. By leveraging appropriate prompting strategies, we enhanced the model’s ability to provide contextually relevant responses.
Secondly, we addressed the issue of data redundancy inherent in EEG signal processing. Through efficient data compression techniques, we optimized the handling of long sequences of EEG data, significantly improving computational efficiency and facilitating real-time emotion computation.
Lastly, we emphasized the importance of patient privacy by ensuring that our model can run locally. We explored model pruning strategies to create a lightweight version of the language model, making it feasible to deploy in environments with limited computational resources while maintaining performance.
However, simply compressing signal data is computationally intensive in many cases, and in complex scenarios, more channel signals are needed. Therefore, if LLMs can generate dense channel signals using only a few channel data and then analyze the emotions of the subjects [63], [64], computational efficiency will be greatly improved.
Our findings indicate that EEG Emotion Copilot has the potential to revolutionize the recognition and treatment of emotions in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will combine lightweight LLMs with multimodal data such as human facial expressions, speech, gestures to achieve greater accuracy in affective computing.

### v2

@@ III-D Model Deployment in EEG Emotion Copilot
Given the various model deployment tools, such as llama.cpp, which can quantify the model based on the specific scenario and convert it into gguf format, it can be directly utilized through ollama, llmstudio, gradio or other tools. This process with gradio ultimately results in the model deployment part illustrated in Fig. 1.
Figure 5: Two fine-tuning strategies. Strategy 1 involves initially training the model on a specific EEG dataset tailored to the current task. Strategy 2 begins with training the model on public datasets, and subsequently fine-tuning it on a specific EEG dataset.
Figure 6: Model’s fine-tuning process on the specific EEG dataset. This fine-tuning process is divided into five steps, each consisting of three epochs. The first four steps utilize 250 Hz data from 55 subjects. Comparing the model’s answers in steps 3 and 4 reveals that the answer in step 3 is richer in content, with no significant difference in training loss between the two steps. This suggests that overtraining lightweight models may be redundant or even detrimental. In step 5, 1000 Hz data from 68 subjects is used for training, employing the pre-trained model from step 4.
Algorithm 2 Pruning Process
1:Input: Initial model $\mathcal{M}$, Parameter group set $\mathcal{P^{\prime}}$, importance function $\mathcal{I}$, contraction strength $\alpha$
2:Output: Pruned model $\mathcal{M}^{\prime}$
3:Calculate group importance:
4:for all $p^{\prime}\in\mathcal{P^{\prime}}$ do
5:     $\mathcal{I}(p^{\prime})\leftarrow\sum_{w\in\mathcal{P^{\prime}}}\|w\|_{2}^{2}$ (See Formula 3, 4, 5, 6)
6:end for
7:    $\mathcal{P}^{\prime\prime}\leftarrow\text{sort}(\mathcal{P^{\prime}},\text{by }\mathcal{I})$
8:Select groups for pruning based on importance:
9:for all $p^{\prime\prime}\in\mathcal{P}^{\prime\prime}$ do
10:     if $\mathcal{I}(p^{\prime\prime})$leq $\theta$ is satisfied then #$\theta$ is the threshold
11:         $\mathcal{P^{\prime\prime}}\leftarrow\mathcal{P^{\prime\prime}}\setminus p^{\prime\prime}$
12:     end if
13:end for
14:return $\mathcal{M}^{\prime}$
III-E Generation of Assisted Electronic Medical Records
Due to the fact that the proposed copilot involves various tasks such as emotion recognition, diagnosis, and treatment plan design, it is sufficient to embed demographic data from different cases under various emotional states, along with the corresponding compressed EEG signals, into the training data. In downstream applications, participant data can be input into the fine-tuned model in a specific format to obtain diagnostic information and treatment plans, as shown in Fig. 7.
Figure 7: Flowchart of assisted electronic medical record generation implemented in EEG Emotion Copilot.
[…]
@@ IV-A Implementation Details
Dataset: We utilize the FACED dataset [29] to validate our copilot. This dataset includes recordings from 123 subjects, each with 32 EEG channels, as they watched 28 video clips, each lasting 30 seconds. These clips encompass nine emotions: anger, disgust, fear, sadness, neutral, amusement, Inspiration, Joy, and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 and then trained on the public datasets with a learning rate of 1e-5, incorporating L1 and L2 regularization terms. The training was conducted for 40 Epoches. During the LoRA fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRAr to 8, LoRAα to 32, and LoRAdropout to 0.5. The fine-tuning process spanned 15 epochs, dividing into 5 steps. Model pruning was executed on a NVIDIA A800 80G, while the fine-tuning were completed on NVIDIA GeForce RTX 4090 GPUs.
@@ IV-B Performance Validation of Varied EEG Signal Length
Regarding the effect of the compressed signal length, we first extracted the EEG signal in the FACED data and the corresponding demography according to the Data Structure in Fig. 1. We implemented two compression methods and an original control group. The question-answering process and the corresponding results are shown in Fig. 8. Although the euclidean distance and cosine similarity of Ans1 and Ans2 are similar to the baseline for the same question, Ans2 is more consistent with the baseline than Ans1, as Ans1 differs from the baseline in terms of emotion recognition. Therefore, with limited resources, it is often more effective to compress the signal to a fixed length using wavelet transforms rather than a segmented approach.
TABLE I: Performance Comparison Of Competing Light-Weight Models
Compression Model Parameters Storage_size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 Avg. RT (s) F1 Avg. RT (s)
[…]
checkpoint1-1 0.15B 0.294G bf16 3 0.080 0.430 0.216 0.420
checkpoint1-2 0.15B 0.294G bf16 3 0.096 0.510 0.324 0.510
checkpoint1-3 0.15B 0.294G bf16 3 0.100 0.540 0.376 0.540
checkpoint1-4 0.15B 0.294G bf16 3 0.093 0.600 0.359 0.610
checkpoint1-5 0.15B 0.294G bf16 3 0.103 0.610 0.340 0.580
base_model-1 [22] 0.50B 0.920G bf16 6 0.143 1.120 0.445 1.220
opt-350m-1 [58] 0.35B 0.647G bf16 15 0.036 3.240 0.037 3.180
LiteLlama 0.46B 0.901G bf16 15 0.041 1.910 0.057 1.900
gpt-sw3-356m 0.47B 1.580G bf16 15 0.035 1.160 0.035 1.180
W Our Model 2
checkpoint2-1 0.15B 0.294G bf16 3 0.118 0.580 0.432 0.590
checkpoint2-2 0.15B 0.294G bf16 3 0.346 0.590 0.980 0.590
checkpoint2-3 0.15B 0.294G bf16 3 0.339 0.550 0.960 0.540
base_model-2 [22] 0.50B 0.920G bf16 6 0.153 1.230 0.450 1.290
opt-350m [58] 0.35B 0.647G bf16 15 0.037 3.960 0.038 4.090
LiteLlama 0.46B 0.901G bf16 15 0.036 2.120 0.036 2.120
gpt-sw3-356m 0.47B 1.580G bf16 15 0.036 1.070 0.036 1.260
[…]
IV-C Performance Comparison with Lightweight Models
To better validate our experimental hypothesis, we combine the strategies outlined in Fig. 6 and Fig. 8. First, we divide the training process based on the data derived from compression method 1, into five steps, with each step trained for 3 epochs. For each step’s checkpoint (i.e., checkpoint1-n), we perform classification of content generation for the nine emotions described in III-A, followed by classification for three emotions (positive, negative, neutral). For comparison, we use Qwen2-0.5B as the base model, training it for 6 epochs under identical experimental conditions to obtain base_model-1, and perform classification, as shown in Table I.
Next, we fine-tuned the models based on the data derived from compression method 2. Specifically, we fine-tuned the pruned model in three steps, with 3, 3, and 3 epochs respectively, and perform classification for both the nine emotions and the three emotions. For consistency, Qwen2-0.5B is also trained for 6 epochs to obtain base_model-2, after which the corresponding classification experiment is conducted, as shown in Table I.
Additionally, we include opt-350m [58], LiteLlama, and gpt-sw3-356m as baseline references. Indeed, the checkpoint1-n model of Model 1 did not achieve ideal results on both tasks compared to the base_model-1. However, considering the F1 scores across multiple epochs of training, the model’s performance on this task gradually improved. In Model 2, the performance of checkpoint2-2 and checkpoint2-3 showed significant improvement compared to models with the similar parameter size, and they also consumed less average response time.
IV-D Performance Comparison with Heavyweight Models
In Table II, we present the results of emotion recognition for Qwen2-1.5B [22], InternLM2.5-1.8B [59], Qwen2.5-3B [60], and Qwen2.5-7B [60], and InternLM2.5-7B [59]. Surprisingly, the results were not as impressive as expected. The reason for this is that the default setting for top-k is 50. But even if top-k is set to 1, the results are still not ideal. For the heavyweight models, their responses are more flexible due to their larger knowledge storage, allowing them to generate a wider range of answers. This also indicates that, despite the very low loss during training, which might suggest model fitting, the actual responses in specific scenarios are dissatisfactory.
Figure 9: t-SNE [61] visualization of K-Means (A-B) and spectral clustering (C-D) results of embeddings based on the EEG signal length=500 (compression method 1) and EEG signal length=50 (compression method 2) in Fig. 8, respectively.
Figure 10: Clustering results of data for compression method 1 and 2 described in Fig. 8 using various models mentioned in Table I and II with K-Means clustering and t-SNE [61] visualization of the text embeddings.
Furthermore, we first use the pre-trained all-MiniLM-L6-v2 model for text representation. Specifically, we combine the “prompt” and “treatment” from each record into a single string and generate the corresponding embedding vectors. These embeddings are then subjected to K-Means clustering and spectral clustering [62] for unsupervised learning, with the number of clusters set to 9. The t-SNE [61] is applied to reduce the dimensionality of the embedding and visualize the results, as shown in Fig. 9.
In addition, we investigate the clustering performance for data of compression method 1 and 2 respectively in Fig. 8 by applying the same procedure to a range of models included in Table I and II. We convert the text into the appropriate input format using the tokenizer, extract the hidden states from the final layer, and apply average pooling to generate fixed-length text embedding vectors. These embeddings are clustered using K-Means, and dimensionality reduction is performed using the t-SNE for visualization, as shown in Fig. 10. Indeed, as illustrated in Fig. 9 and 10, the visualization suggests that corpora of two different lengths are not easily clustered. Specifically, compared to the data of Model 1, the data of Model 2 exhibits a higher degree of aggregation, which directly corresponds to a reduction in computational load.
Figure 11: Performance comparison as to F1 and Avg. RT of nine and three emotion classification across steps of two fine-tuning strategies.
IV-E Model Fine-tuning Strategy Comparison After Pruning
To evaluate the two fine-tuning strategies outlined in Fig. 5, we adopted compression method 2 show in Fig. 8 for comparison. According to the results from Table I and Fig. 11, we found that Strategy 1 demonstrated better performance in Table I than Strategy 2 did in the lightweight models fine-tuning process.
TABLE II: Performance comparison Of Heavy-Weight Models
Compression Model Parameters Storage Size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 (Top-K=50) Avg. RT (s) F1 (Top-K=50) Avg. RT (s)
[…]
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 5.870 0.035 5.910
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.111 4.260 0.041 10.960
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.051 1.410 0.152 2.100
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.082 10.090 0.039 10.090
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.074 3.730 0.328 3.510
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 6.320 0.041 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.133 3.680 0.114 3.760
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.036 0.020 0.038 0.240
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.082 10.260 0.042 7.300
Compression Model Parameters Storage Size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 (Top-K=1) Avg. RT (s) F1 (Top-K=1) Avg. RT (s)
[…]
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 7.100 0.036 5.830
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.080 3.810 0.342 4.500
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.040 0.690 0.036 0.870
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.090 8.950 0.040 9.390
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.081 3.780 0.292 4.170
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 7.210 0.036 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.036 6.320 0.038 4.470
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.051 4.870 0.110 4.920
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.081 11.640 0.041 9.790
IV-F Assisted Electronic Medical Record Generation with Our Copilot
@@
Fig. 12 illustrates an example of using our EEG Emotion Copilot to create an electronic medical record. Initially, the starting sentence of the prompt is provided, followed by the addition of demographic information of the subjects, including age, gender, and certain facial features. Finally, the preprocessed and compressed signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot. The EEG Emotion Copilot first analyzes the data to generate an answer, which is subsequently transformed into the record through format conversion.
V Discussion
V-A Data Structure of Prompt
We developed a comprehensive human-based dataset that integrates both public knowledge datasets and specialized emotion datasets, facilitating the construction of a robust data structure to drive the development of LLMs. By employing effective prompting strategies, we enhanced the model’s ability to generate contextually relevant responses. Furthermore, while the performance of the models may not be exceptionally high, as illustrated in Table I and II, language remains more challenging to cluster unsupervised compared to other data types [63]. Nonetheless, by utilizing a pretrained sentence transformer model of an appropriate size, the model is still able to capture the underlying emotional features of the language to some extent as demonstrated in the spectral clustering shown in Fig. 9, despite potential issues with feature imbalance.
V-B Data Redundancy in EEG Signal Processing
We addressed the issue of data redundancy inherent in EEG signal processing. Through efficient data compression techniques, we optimized the handling of long sequences of EEG data, significantly improving computational efficiency and facilitating real-time emotion computation. However, simply compressing signal data is computationally intensive in many cases, and in complex scenarios, more channel signals are needed. Therefore, if LLMs can generate dense channel signals using only a few channel data and then analyze the emotions of the subjects [64], [65], computational efficiency will be greatly improved.
V-C Patient Privacy and Model Efficiency
We emphasized the importance of patient privacy by ensuring that our model can run locally. We explored model pruning strategies to create a lightweight version of the language model, making it feasible to deploy in environments with limited computational resources while maintaining performance.
V-D Performance of Lightweight Models
We performed model pruning and compared the performance of the models obtained through the two fine-tuning methods illustrated in Fig. 5. It was found that the results obtained using Strategy 1 for fine-tuning were superior. Furthermore, as shown in Table I and II, the lightweight pruned model achieved through multi-step fine-tuning demonstrated better performance compared to other models with larger parameter sizes.
V-E Hallucination of Heavyweight Models
While a low training loss suggests a good fit to the training data, it does not necessarily reflect the model’s true performance, as the generated responses may still fail to meet expectations or even diverge significantly [66]. This issue is particularly evident in 7B and larger models, where the hallucination problem becomes more pronounced in specialized domains. These findings also suggest that lightweight models may be better suited for such domain-specific generation tasks.
@@ V-E Hallucination of Heavyweight Models
VI Conclusion
In this paper, we present EEG Emotion Copilot, a lightweight and locally-running LLM designed for emotional recognition using EEG signals and the generation of corresponding treatment plans for assisted electronic medical records. Our findings indicate that EEG Emotion Copilot has the potential to revolutionize the recognition and treatment of emotions in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will combine lightweight LLMs with multimodal data such as human facial expressions, speech, gestures to achieve greater accuracy in affective computing.